OpenStudy (dtan5457):

How would I simplify sin2(theta)/2cos(theta)?

OpenStudy (dtan5457):

@freckles

OpenStudy (dtan5457):

@misty1212

OpenStudy (dtan5457):

@rational

zepdrix (zepdrix):

\[\Large\rm \frac{\color{orangered}{\sin (2\theta)}}{2\cos\theta}\]Hey David :) Apply your Sine Double Angle Formula to the numerator.

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{\sin(2\theta)=2\sin \theta \cos \theta}\]Remember that one?

OpenStudy (dtan5457):

Yeah

OpenStudy (dtan5457):

I got that part and the bottom Im not sure which one to use

zepdrix (zepdrix):

No no, leave the bottom alone. That's just a simple cosine down there, ya? Change the top, then look for cancellations.

OpenStudy (dtan5457):

I'm getting just sinA, is that right?

zepdrix (zepdrix):

Yay good job \c:/ sin(theta) appears to be all we're left with!

OpenStudy (dtan5457):

ty